# Discrete Differential Geometry for $C^{1,1}$ Hyperbolic Surfaces of Non-Constant Curvature

### v3

@@
Abstract. We develop a discrete differential geometry for surfaces of non-constant negative curvature, which can be used to model various phenomena from the growth of flower petals to marine invertebrate swimming. Specifically, we derive and numerically integrate a version of the classical Lelieuvre formulas that apply to immersions of $C^{1,1}$ hyperbolic surfaces of non-constant curvature. In contrast to the constant curvature case, these formulas do not provide an explicit method for constructing an immersion but rather describe an immersion via an implicit set of equations. We propose an iterative method for resolving these equations. Because we are interested in scenarios where the curvature is a function of the intrinsic material coordinates, in particular, on the geodesic distance from an origin, we suggest a fast marching method for computing geodesic distance on manifolds. We apply our methods to generate surfaces of non-constant curvature and demonstrate how one can introduce branch points to account for the sort of multi-generational buckling and subwrinkling observed in many applications.
@@ 1 Introduction
Thin elastic sheets which bend, buckle, and wrinkle arise in a variety of physical phenomena, both natural and man-made. Some applications where these appear are biological processes like leafy growth or marine flatworm locomotion [HWQ+18, NV01, YSS+21], the compression of steel during automotive collisions [CCY+04, MAA22, MFS11], the irregular crenellations at the edges of torn plastic [SRS07], and photo- and thermo-sensitive hydrogels [LSH13, PPK23].
@@
Over the last decade, the theory of non-Euclidean elastic plates has been developed to model such phenomena [ESK09, LP11]. A key underlying assumption is that the configurations of these “plates” that we observe are minimizers of an elastic energy functional. The goal is then to create immersions of a finite hyperbolic surfaces into $\mathbb{R}^{3}$ which minimize elastic energy, and the analysis of these immersions is central to some recent work devoted to generating these surfaces [GSSV16, GV13, Ven19]. For an overview of the the calculus of variations for thin elastic sheets, we direct the reader to [Lew23].
@@
In this manuscript, we will be concerned with the modeling of 2-dimensional Riemannian surface $(\Omega,g)$ of non-constant negative curvature realized as an immersion into $\mathbb{R}^{3}$. Mathematically, this amounts to solving a nonlinear PDE, the prescribed Gauss curvature equation [Sto89] for an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$. This equation is a fully nonlinear hyperbolic equation if the prescribed curvature is negative.
@@
In particular, motivated by the $C^{1}$ immersion theorem of Nash-Kuiper [Kui55, Nas54] and the nonexistence of $C^{2}$ immersions for complete (unbounded) hyperbolic surfaces proven by Efimov [Efi62, Mil72], Shearman and Venkataramani [SV21] developed the theory of distributed branch points for negatively curved surfaces, which allows immersions whose smoothness is strictly between $C^{1}$ and $C^{2}$. They argue that such immersions can have significantly lower elastic energy than $C^{2}$ immersions of finite pieces of hyperbolic surfaces, especially as these pieces become large, and can successfully capture the sort of multigenerational buckling and subwrinkling that appear in nature and are not captured by smoother immersions. They also develop discrete differential geometry (DDG) for such surfaces, and suggest a method for constructing discrete immersions by surgery, i.e. recursively patching together “sectors” — bounded hyperbolic surfaces that admit a pair of asymptotic coordinates [Sto89] and have data specified along two transversely intersecting boundary curves, i.e each sector is a solution to a Goursat problem [HW52].
@@
In general, one cannot expect to obtain solutions for a nonlinear PDE as a piecewise function obtained from local solutions. However, the prescribed Gauss curvature equation for immersions of hyperbolic surfaces has a special structure — it is the compatibility condition [BS08] for the Lelieuvre equations [Eis97], a pair of ODEs defining the immersion $\mathbf{r}$ along asymptotic curves [HW51, Sto89] which are the characteristic curves for the underlying hyperbolic PDE. The key insight in [SV21] was that if two solutions agree along an asymptotic curve, then they can be patched together along this curve to obtain a piecewise defined weak solution for an immersion with a prescribed Gauss curvature. This naturally leads to the question of obtaining solutions on sectors, i.e. regions bounded by a pair of asymptotic curves. Each sector is then obtained by solving a Goursat problem, i.e. a second order hyperbolic PDE with data specified along two characteristic curves [HW52], which, with appropriate choices for the boundary data, can be patched together with other sectors. In particular, this procedure can be carried out using sectors where one or both boundaries are straight lines in $\mathbb{R}^{3}$. Hyperbolic surfaces that contain two intersecting straight lines were first obtained by Amsler [Ams55] and their discrete analogs have also been studied [BMS02]. Each of the four quadrants defined by these intersecting lines is an example of an Amsler sector [SV21]. In addition, the surgery procedure also uses Pseudo-Amsler sectors – sectors where one of the boundary curves is a straight line, but the other boundary is a prescribed space curve in $\mathbb{R}^{3}$ [SV21].
@@
All the work in [SV21]—as well as that in [GSSV16, GV13]—applies to surfaces of constant negative curvature. In this work, we extend these methods to surfaces of non-constant negative curvature. In physical scenarios, this could account for thin sheets that are relatively flat in some places but become increasingly curvy in others. We are specifically interested in scenarios where curvature depends on the geodesic distance from some “center” point—which could accurately model objects like leaves or flower petals that become frilly only near the edges—though the methods described could be applied more generally. A key tool for us will be the Lelieuvre formulas [Eis97, Chap 6]. These formulas establish a relationship between an immersion of the surface and the normal vector along the surface and provide a relatively simple way to generate hyperbolic surfaces with prescribed negative curvature by evolving the surface “outward” from consistently defined boundary data. However, in the case of non-constant curvature, this introduces some difficulty in synthesizing the different coordinate systems used to describe the surface. To briefly explain, a thin elastic sheet which we are describing mathematically as a two-dimensional hyperbolic surface will have material coordinates that can be thought of as intrinsic to the object itself. Alternatively, we can describe the sheet via immersion into Euclidean space, whereupon we introduce a set of three-dimensional coordinates as well as asymptotic coordinates which are specific to the immersion and thus extrinsic to the object. In analogy to fluid dynamics, the former could be seen as the Lagrangian frame and the latter the Eulerian frame (see [RKN21] for a discussion of this distinction in fluid dynamics, both theoretically and computationally). The Lelieuvre formulas are expressed in asymptotic coordinates but depend on the prescribed curvature which is intrinsic and should thus be expressed in terms of intrinsic quantities (such as the material coordinates or geodesic distance to a reference point). The connection between these intrinsic and extrinsic coordinates is the immersion itself, and this creates a dependency loop: the Lelieuvre equations are used to construct the immersion of the surface into Euclidean space, but depend on intrinsic information which can only be resolved once the immersion is already fully known. Much of the work contained in the manuscript is devoted to overcoming this difficulty.
@@
The organization of this document is as follows. In section 2, we present a self-contained derivation of the Lelieuvre formulas for a surface of non-constant negative curvature. In section 3, we develop the discrete analog to the continuous Lelieuvre formulas. In section 4, we describe a fast marching method for computing geodesic distance on triangulated manifolds, which will allow us to generate surfaces with prescribed negative curvature depending on the geodesic distance. In section 5, we apply our methods to generate Amsler surfaces with non-constant negative curvature. Finally, in section 6, we describe how the method of “surgery” developed by [SV21] can be adapted to our scenario, and thus generate $C^{1,1}$ surfaces with branch points that allow for buckling and subwrinkling.
@@ 2 Lelieuvre Formulas for Surfaces with Non-Constant Negative Curvature
In this section, we present a short and self-contained derivation of the Lelieuvre formulas [RS02], a set of differential equations describing the geometry of negatively curved surfaces. Our thin elastic object is modeled by a 2-dimensional Riemannian manifold $(\Omega,g)$ representing its mid-surface. The metric tensor $g$ determines the intrinsic geometry on $\Omega$ [Sto89], and is called the target metric in the context of non-Euclidean elasticity [ESK09].
Supposing our surface $\Omega$ has a pair of reference coordinates $(x,y)$, we begin with an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ of our surface into Euclidean space. The immersion induces a metric $I$ on $\Omega$, called the first fundamental form [Sto89], by pulling back the standard Euclidean metric on $\mathbb{R}^{3}$, i.e. $I=d\mathbf{r}\cdot d\mathbf{r}$. The difference $I-g$ between the induced and the target metrics on $\Omega$ is a measure of the stretching of the sheet. In the thin limit, sufficiently regular elastic objects are essentially unstretchable [LP11], so the allowed configurations satisfy the isometry constraint $d\mathbf{r}\cdot d\mathbf{r}=g$.
@@
The map $\mathbf{N}:\Omega\to S^{2}$ given by $p\mapsto\mathbf{N}(p)$ is called the Gauss normal map [Sto89]. Fig. 2 shows a schematic representation of these mappings.
@@
Note that for a $C^{k}$ surface $\mathbf{r}(x,y)$ the normal map is only $C^{k-1}$. If the normal map is differentiable at a point $p$, by differentiating $\left|\mathbf{N}\right|=1$, we see $\mathbf{N}\cdot\mathbf{N}_{x}=\mathbf{N}\cdot\mathbf{N}_{y}=0$, which implies that $\mathbf{N}_{x}$ and $\mathbf{N}_{y}$ also lie in the tangent plane $T_{p}$. The shape operator at the point $p$ is the unique linear operator $S_{p}:T_{p}\to T_{p}$ such that $S_{p}(\mathbf{r}_{x}(p))=\mathbf{N}_{x}(p)$ and $S_{p}(\mathbf{r}_{y}(p))=\mathbf{N}_{y}(p)$ [Sto89]. The shape operator provides information regarding local curvature of the immersed surface $\mathbf{r}(\Omega)$ near $p$. The determinant of the shape operator defines the Gaussian curvature extrinsically: $K(p;\mathbf{r}):=\text{det}(S_{p})$. From this, we see
@@
or $\left|K(p;\mathbf{r})\right|=\frac{\left|\mathbf{N}_{x}\times\mathbf{N}_{y}\right|}{\left|\mathbf{r}_{x}\times\mathbf{r}_{y}\right|}$ with the sign determined by the relative orientations of $\mathbf{N}$ and $\mathbf{r}_{x}\times\mathbf{r}_{y}$. This states that the Gauss curvature is the local magnification of an infinitesimal area element on the surface under the normal map $\mathbf{N}:\Omega\to S^{2}$ given by $p\mapsto\mathbf{N}(p)$. We note that a priori $K(p;\mathbf{r})$ may also depend on the immersion $\mathbf{r}$. However, as a consequence of the Theorema Egregium, Gauss curvature is intrinsic [Sto89] which implies that if $d\mathbf{r}\cdot d\mathbf{r}=g$, then $K(p;\mathbf{r})$ is determined by $g$ and independent of the particular immersion $\mathbf{r}$. This yields the prescribed Gauss curvature equation
@@
For a broader context, equation (3) is an example of a Monge-Ampère equation, that is hyperbolic (resp. elliptic) if $K_{g}<0$ (resp. $K_{g}>0$). In contrast to the case of convex solutions of elliptic Monge-Ampere equations, that have a well-developed regularity theory [Pog64, Caf90], the regularity theory for the hyperbolic case, and even the question of the appropriate definition of the notion of a weak solution, (when considered on sets of various different codimensions in spaces of various different dimensions) is very much an active area of research [CI23, CS19, CSJ25, Lew25a, Lew25b, LP17]. In this work, we will be primarily concerned with immersions $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ that are $C^{1,1}$ and piecewise smooth, but not necessarily globally $C^{2}$. Despite the lack of two derivatives for $\mathbf{r}$ it is straightforward to interpret such $C^{1,1}$ immersions as weak solutions of (3) by multiplying both sides of the equation by compactly supported smooth functions and integrating over the domain $\Omega$. Indeed, we only need a.e. defined measurable functions $\text{det}(S_{p})$ and $K_{g}(p)$ for this procedure to make sense. In our setting $K_{g}$ is a “fixed” smooth function, i.e. independent of the immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$. For a $C^{1,1}$ immersion $\mathbf{r}$, the normal map $\mathbf{N}:\Omega\to S^{2}$ is Lipschitz, and Rademacher’s theorem [Hei05, §3] implies that $\mathbf{N}$ is a.e. differentiable, so the shape operator $S_{p}$ is defined a.e.
If an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ is smooth in a neighborhood of a point $p$ and the Gauss curvature is negative at a point $p$, then the shape operator $S_{p}$ is well defined at $p$ and has one negative and one positive eigenvalue, and thus there are two linearly independent vectors in $T_{p}$ each satisfying $\mathbf{v}\cdot S_{p}(\mathbf{v})=0$. These vectors define the asymptotic directions at $p$: the directions along which there is zero normal curvature [Sto89]. Assuming the Gaussian curvature is negative everywhere and the vectors defining these asymptotic directions vary smoothly with $p$, we can extend them to integral curves which give asymptotic coordinates on the manifold [HW51, Sto89]. We denote the asymptotic coordinates by $(u,v)$. These coordinates are not intrinsic to the surface, i.e. tied to material points, but rather they depend on the particular immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$.
@@
The first and second terms in each expression in (6) are normal to $\mathbf{N}$, so for the equation to hold, the last terms must be equal which shows that $a(p)=-b(p)$. From (2), we see $a(p)b(p)=K(p)^{-1}$ so that
@@
and similarly in the equation for $\mathbf{r}_{v}$. Finally, defining the scaled normal $\boldsymbol{\nu}=(-K(p)^{-1/4})\mathbf{N}$ [Sch17], we arrive at
@@
The formulas in (9) and (10) are a version of the Lelieuvre formulas [Eis97] with non-constant curvature. For pseudospherical surfaces—those with constant negative curvature—one can choose units of length so that the curvature is $K(p)=-1$, which gives the classical Leliuvre formulas (and compatibility condition):
@@
In this manuscript, we consider curvature functions which are nonconstant, but are negative and bounded away from zero, whereupon by changing length scales, we can assume that $K(p)\leq-1$ without loss of generality. This will be a tacit assumption throughout.

### v4

@@
Abstract. We develop a discrete differential geometry for surfaces of non-constant negative curvature, which can be used to model various phenomena from the growth of flower petals to marine invertebrate swimming. Specifically, we derive and numerically integrate a version of the classical Lelieuvre formulas that apply to immersions of $C^{1,1}$ hyperbolic surfaces of non-constant curvature. In contrast to the constant curvature case, these formulas do not provide an explicit method for constructing an immersion but rather describe an immersion via an implicit set of equations. We propose an iterative method for resolving these equations. Because we are interested in scenarios where the curvature is a function of the intrinsic material coordinates, in particular, on the geodesic distance from an origin or from an edge, we suggest a fast marching method for computing geodesic distance on manifolds. We apply our methods to generate surfaces of non-constant curvature and demonstrate how one can introduce branch points to account for the multi-generational buckling and subwrinkling observed in many applications.
@@ 1 Introduction
Thin elastic sheets that bend, buckle, and wrinkle arise in a variety of physical phenomena, both natural and man-made. Some applications where these appear are biological processes like leafy growth or marine flatworm locomotion [26, 42, 58], the compression of steel during automotive collisions [15, 38, 40], the irregular crenellations at the edges of torn plastic [51, 52], and photo- and thermo-sensitive hydrogels [37, 46].
@@
Over the last decade, the theory of non-Euclidean elastic plates has been developed to model such phenomena [18, 34]. A key underlying assumption is that the observed configurations of these plates are minimizers of an elastic energy functional [18]. The goal is then to create immersions of bounded hyperbolic surfaces into $\mathbb{R}^{3}$ that minimize elastic energy, and the analysis of these immersions is central to some recent work devoted to generating these surfaces [20, 21, 56]. For an overview of the calculus of variations for thin elastic sheets, we direct the reader to [31].
@@
In the limit of vanishing thickness, the sheet becomes essentially inextensible and the energy minimization problem enforces a hard isometry constraint on the immersion [34]. In this manuscript, we will be concerned with modeling 2-dimensional Riemannian surfaces $(\Omega,g)$ of non-constant negative curvature realized as isometric immersions into $\mathbb{R}^{3}$. Mathematically, this amounts to solving a nonlinear PDE, the prescribed Gauss curvature equation [54], for an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$. This equation is a fully nonlinear hyperbolic Monge-Ampère equation if the prescribed curvature is negative.
A separate, yet related, line of inquiry in the field of geometric mechanics concerns the application of the Monge-Ampère equation to prestrained plates with positive curvature, which are governed by the elliptic Monge-Ampère equation. While the analytic theory for this equation is well-established, the development of corresponding computational methods is a more recent endeavor. Difficulties arise from several sources. Beyond the equation’s inherent nonlinearity, weak solutions are not based on standard variational principles but on concepts such as viscosity solutions, which can be challenging to interpret discretely [44]. Furthermore, the global convexity constraint, often required for a well-posed theory, is difficult to enforce in a numerical setting. Consequently, the construction of convergent numerical schemes remains an active area of research, even for this well-studied, elliptic class of Monge-Ampère equations [43].
In contrast, even the analytic theories for parabolic and hyperbolic Monge-Ampère equations are very much under active development. On the numerical side, there have been several recent developments in the simulation of the parabolic Monge-Ampère equation [2], which arises in applications such as curved crease origami in flat sheets [36]. The numerical solution of this equation is complicated by the presence of curved creases, for which many standard methods fail to converge [4]. This highlights the distinct challenges posed by different classes of Monge-Ampère equations and the motivation for developing specialized numerical techniques (See [10, 9] and references therein).
@@
In this paper, we develop methods for the hyperbolic setting. In particular, motivated by the $C^{1}$ immersion theorem of Nash-Kuiper [30, 41] and the nonexistence of $C^{2}$ immersions for complete (unbounded) hyperbolic surfaces proven by Efimov [17, 39], Shearman and Venkataramani [53] developed the theory of distributed branch points for negatively curved surfaces, which allows immersions whose smoothness is strictly between $C^{1}$ and $C^{2}$. They argue that such immersions can have significantly lower elastic energy than $C^{2}$ immersions of finite pieces of hyperbolic surfaces, especially as these pieces become large, and can successfully capture the sort of multigenerational buckling and subwrinkling that appear in nature and are not captured by smoother immersions. They also develop discrete differential geometry (DDG) for such surfaces, and suggest a method for constructing discrete immersions by surgery, i.e. recursively patching together “sectors” — bounded hyperbolic surfaces that admit a pair of asymptotic coordinates [54] and have data specified along two transversely intersecting boundary curves [23].
It is worth emphasizing that the methods developed in [53], and extended in this manuscript, are intrinsically discrete rather than being discrete approximations of a continuous formulation. Our approach borrows extensively from the concepts of Discrete Differential Geometry (DDG), which aims to construct discrete theories that run parallel to the smooth theory while preserving its essential geometric and structural properties. This philosophy guides our development of the discrete Lelieuvre formulas in Section 3.
@@
In general, one cannot expect to obtain solutions for a nonlinear PDE as a piecewise function obtained from local solutions. However, the prescribed Gauss curvature equation for immersions of hyperbolic surfaces has a special structure — it is the compatibility condition [8] for the Lelieuvre equations [19], a pair of ODEs defining the immersion $\mathbf{r}$ along asymptotic curves [22, 54] which are the characteristic curves for the underlying hyperbolic PDE. The key insight in [53] was that if two solutions agree along an asymptotic curve, then they can be patched together along this curve to obtain a piecewise defined weak solution for an immersion with a prescribed Gauss curvature. This naturally leads to the question of obtaining solutions on sectors, i.e. regions bounded by a pair of asymptotic curves. Each sector is then obtained by solving a Goursat problem, i.e. a second order hyperbolic PDE with data specified along two characteristic curves [23], which, with appropriate choices for the boundary data, can be patched together with other sectors. In particular, this procedure can be carried out using sectors where one or both boundaries are straight lines in $\mathbb{R}^{3}$. Hyperbolic surfaces that contain two intersecting straight lines were first obtained by Amsler [1] and their discrete analogs have also been studied [5]. Each of the four quadrants defined by these intersecting lines is an example of an Amsler sector [53]. In addition, the surgery procedure also uses Pseudo-Amsler sectors – sectors where one of the boundary curves is a straight line, but the other boundary is a prescribed space curve in $\mathbb{R}^{3}$ [53].
@@
All the work in [53]—as well as that in [20, 21]—applies to surfaces of constant negative curvature. In this work, we extend these methods to surfaces of non-constant negative curvature. In physical scenarios, this could account for thin sheets that are relatively flat in some places but become increasingly curvy in others. We are specifically interested in scenarios where curvature depends on the geodesic distance from some “center” or an “edge”—which could accurately model objects like leaves or flower petals that become frilly only near the edges—though the methods described could be applied more generally. A key tool for us will be the Lelieuvre formulas [19, Chap 6]. These formulas establish a relationship between an immersion of the surface and the normal vector along the surface and provide a relatively simple way to generate hyperbolic surfaces with prescribed negative curvature by evolving the surface “outward” from consistently defined boundary data. However, in the case of non-constant curvature, this introduces some difficulty in synthesizing the different coordinate systems used to describe the surface. The connection between these intrinsic and extrinsic coordinates is the immersion itself. This creates a dependency loop: the Lelieuvre formulas, which are expressed in extrinsic asymptotic coordinates, require the curvature as an input; however, the curvature is defined in terms of intrinsic material coordinates. Resolving this connection—constructing an immersion that depends on information that can only be fully known once the immersion itself is constructed—is a central challenge we overcome in this work.
@@
The organization of this paper is as follows. In section 2, we present a self-contained derivation of the Lelieuvre formulas for a surface of non-constant negative curvature. In section 3, we develop the discrete analog to the continuous Lelieuvre formulas. In section 4, we describe a fast marching method for computing geodesic distance on triangulated manifolds, which will allow us to generate surfaces with prescribed negative curvature depending on the geodesic distance. In section 5, we apply our methods to generate Amsler surfaces with non-constant negative curvature. In section 6, we describe how the method of “surgery” developed by [53] can be adapted to our scenario, and thus generate $C^{1,1}$ surfaces with branch points that allow for buckling and subwrinkling. While sections 5 and 6 utilize a disk-like geometry, in section 7, we present a further application of our methods in a geometry which models the edge of a torn plastic strip as studied in [51, 52]. Finally, in appendix A, we give a proof that our iterative algorithm for generating a discrete surface of prescribed, non-constant, negative Gauss curvature does indeed converge.
@@ 2 Lelieuvre Formulas for Surfaces with Non-Constant Negative Curvature
In this section, we present a short and self-contained derivation of the Lelieuvre formulas [19, Chap 6], a set of differential equations describing the geometry of negatively curved surfaces. Our thin elastic object is modeled by a 2-dimensional Riemannian manifold $(\Omega,g)$ representing its mid-surface. The metric tensor $g$ determines the intrinsic geometry on $\Omega$ [54], and is called the target metric in the context of non-Euclidean elasticity [18].
Supposing our surface $\Omega$ has a pair of reference coordinates $(x,y)$, we begin with an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ of our surface into Euclidean space. The immersion induces a metric $I$ on $\Omega$, called the first fundamental form [54], by pulling back the standard Euclidean metric on $\mathbb{R}^{3}$, i.e. $I=d\mathbf{r}\cdot d\mathbf{r}$. The difference IgI-g between the induced and the target metrics on $\Omega$ is a measure of the stretching of the sheet. In the thin limit, sufficiently regular elastic objects are essentially unstretchable [34], so the allowed configurations satisfy the isometry constraint $d\mathbf{r}\cdot d\mathbf{r}=g$.
@@
The map $\mathbf{N}:\Omega\to S^{2}$ given by $p\mapsto\mathbf{N}(p)$ is called the Gauss normal map [54]. Fig. 2 shows a schematic representation of these mappings.
@@
Note that for a $C^{k}$ surface $\mathbf{r}(x,y)$ the normal map is only $C^{k-1}$. If the normal map is differentiable at a point $p$, by differentiating $\left|\mathbf{N}\right|=1$, we see $\mathbf{N}\cdot\mathbf{N}_{x}=\mathbf{N}\cdot\mathbf{N}_{y}=0$, which implies that $\mathbf{N}_{x}$ and $\mathbf{N}_{y}$ also lie in the tangent plane $T_{p}$. The shape operator at the point $p$ is the unique linear operator $S_{p}:T_{p}\to T_{p}$ such that $S_{p}(\mathbf{r}_{x}(p))=\mathbf{N}_{x}(p)$ and $S_{p}(\mathbf{r}_{y}(p))=\mathbf{N}_{y}(p)$ [54]. The shape operator provides information regarding local curvature of the immersed surface $\mathbf{r}(\Omega)$ near $p$. The determinant of the shape operator defines the Gaussian curvature extrinsically: $K(p;\mathbf{r}):=\text{det}(S_{p})$. From this, we see
@@
or $\left|K(p;\mathbf{r})\right|=\frac{\left|\mathbf{N}_{x}\times\mathbf{N}_{y}\right|}{\left|\mathbf{r}_{x}\times\mathbf{r}_{y}\right|}$ with the sign determined by the relative orientations of $\mathbf{N}$ and $\mathbf{r}_{x}\times\mathbf{r}_{y}$. This states that the Gauss curvature is the local magnification of an infinitesimal area element on the surface under the normal map $\mathbf{N}:\Omega\to S^{2}$ given by $p\mapsto\mathbf{N}(p)$. We note that a priori $K(p;\mathbf{r})$ may also depend on the immersion $\mathbf{r}$. However, as a consequence of the Theorema Egregium, Gauss curvature is intrinsic [54] which implies that if $d\mathbf{r}\cdot d\mathbf{r}=g$, then $K(p;\mathbf{r})$ is determined by $g$ and independent of the particular immersion $\mathbf{r}$. This yields the prescribed Gauss curvature equation
@@
For a broader context, equation (3) is an example of a Monge-Ampère equation, that is hyperbolic (resp. elliptic) if $K_{g}<0$ (resp. $K_{g}>0$). In contrast to the case of convex solutions of elliptic Monge-Ampere equations, that have a well-developed regularity theory [45, 11], the regularity theory for the hyperbolic case, and even the question of the appropriate definition of the notion of a weak solution, (when considered on sets of various different codimensions in spaces of various different dimensions) is very much an active area of research [12, 13, 14, 32, 33, 35]. In this work, we will be primarily concerned with immersions $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ that are $C^{1,1}$ and piecewise smooth, but not necessarily globally $C^{2}$. Despite the lack of two derivatives for $\mathbf{r}$ it is straightforward to interpret such $C^{1,1}$ immersions as weak solutions of (3) by multiplying both sides of the equation by compactly supported smooth functions and integrating over the domain $\Omega$. Indeed, we only need a.e. defined measurable functions $\text{det}(S_{p})$ and $K_{g}(p)$ for this procedure to make sense. In our setting $K_{g}$ is a “fixed” smooth function, i.e. independent of the immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$. For a $C^{1,1}$ immersion $\mathbf{r}$, the normal map $\mathbf{N}:\Omega\to S^{2}$ is Lipschitz, and Rademacher’s theorem [25, §3] implies that $\mathbf{N}$ is a.e. differentiable, so the shape operator $S_{p}$ is defined a.e.
If an immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$ is smooth in a neighborhood of a point $p$ and the Gauss curvature is negative at a point $p$, then the shape operator $S_{p}$ is well defined at $p$ and has one negative and one positive eigenvalue, and thus there are two linearly independent vectors in $T_{p}$ each satisfying $\mathbf{v}\cdot S_{p}(\mathbf{v})=0$. These vectors define the asymptotic directions at $p$: the directions along which there is zero normal curvature [54]. Assuming the Gaussian curvature is negative everywhere and the vectors defining these asymptotic directions vary smoothly with $p$, we can extend them to integral curves which give asymptotic coordinates on the manifold [22, 54]. We denote the asymptotic coordinates by $(u,v)$. These coordinates are not intrinsic to the surface, i.e. tied to material points, but rather they depend on the particular immersion $\mathbf{r}:\Omega\to\mathbb{R}^{3}$.
@@
The first and second terms in each expression in (6) are normal to $\mathbf{N}$, so for the equation to hold, the last terms must be equal which shows that a(p)=b(p)a(p)=-b(p). From (2), we see $a(p)b(p)=K(p)^{-1}$ so that
@@
and similarly in the equation for $\mathbf{r}_{v}$. Finally, defining the scaled normal $\boldsymbol{\nu}=(-K(p)^{-1/4})\mathbf{N}$ [48], we arrive at
@@
The formulas in (9) and (10) are a version of the Lelieuvre formulas [19] with non-constant curvature. For pseudospherical surfaces—those with constant negative curvature—one can choose units of length so that the curvature is K(p)=1K(p)=-1, which gives the classical Leliuvre formulas (and compatibility condition):
@@
In this manuscript, we consider curvature functions which are nonconstant, but are negative and bounded away from zero.